Big Box Store (BBS) has an annual rate of 4% of all sales being returned. In a recent sample of thirty randomly selected sales the number of returns was five. BBS is concerned about the event, and your advice is solicited. (8) (2.5 points) What is the probability that the number of returns is not equal to 4 in a random sample of 30 sales?
Answer #1

solution No of wage = (30 (4) = 27405 and prob (no fretroren 9) = (309) 10.09,* (0 763304 = 3044 (0-09) 4 6 4626 =0.024273 vithank you